# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v1

@@ B.3 Dependence on Latent Layer of Study
Figure 14: This network shows the PPP method applied over networks of $3$ hidden layer block, $18$ neurons and using input-output normalisation as described in Apps. A and D. PPP was applied to the $0^{\text{th}}$ hidden layer. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. One can observe that representations are absent in anisotropic rows 1, 2, and 4, likely a result of anti-alignment, whereas some axis-aligned structure is present in rows 3 and 5. Isotropic rows demonstrate larger, more continuous representational structures, but also demonstrate clustering such that they are not distributed smoothly over the representation space.
@@
Overall, in many networks tested, anisotropic structure appeared progressively clearer and denser in later latent layers, whilst isotropic networks tended to produce progressively more continuous distributions in later latent layers. This seemed to be a trend across many of the networks studied, but also sporadic network-to-network differences occurred. This may be due to the increased number of an/isotropies in the forward pass or indirectly through the backwards pass, yet further study may elucidate the exact reason why progressively clearer structure may emerge in later layers.
@@ B.4 Results for Leaky-ReLU
In this section, the standard Leaky-ReLU, shown in Eqn. 6, is compared with its isotropic counterpart, shown in Eqn. 7 [2].
@@
The permutation symmetry of standard Leaky-ReLU subtly differs from that of standard tanh. Standard tanh is equivariant under both permutation and component sign-flips, termed a hyperoctahedral symmetry $B_{n}$, whereas standard Leaky-ReLU is only equivariant under permutation ($S_{n}$). Differences in quantised structure may be indicative of such differences between symmetry groups obeyed by the functional forms.
@@
Figure 18: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with $3$ hidden layer blocks and width of $18$ neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. All five of the top five rows demonstrate clear anti-aligned quantisation in representations, predominantly in a single anti-aligned axis. These emerge early in training and become progressively more (approximately) discrete. This contrasts with the isotropic Leaky-ReLU examples, which approach a more continuous, evenly distributed embedding distribution, consistent with previous observations. Additionally, at relatively smaller magnitudes, the Leaky-ReLU results consistently show some deviation from the discrete-like rays; this remains unexplained.
Figure 19: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with $2$ hidden layer blocks and width of $18$ neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. These results are highly consistent with Fig. 18, yet the 3rd row also shows the addition of axis-aligned structure.
@@
In particular, the $0^{\text{th}}$ layer standard Leaky-ReLU results always produced two narrow beams arranged antipodally, along a single diagonal. This anti-aligned structure occurs in most cases, and the standard Leaky-ReLU only rarely displays an axis-aligned structure. This single-diagonal may be indicative of the $S_{n}$ group, which does not include sign flips, unlike tanh’s $B_{n}$ group. Therefore, only a single diagonal may appear populated, as opposed to other diagonal directions, which could result from the various mirrors applicable in the $B_{n}$ group. This nuanced difference between demonstrations appears to closely follow from the differences between symmetry groups. These differing representations can be considered to support the influence of different symmetry groups on representations, with nuances following in a taxonomic fashion from the group structures.
@@
Particularly, Leaky-ReLU effectively uses the condition that the activation is scaled by $\alpha$ if any component of the vector, decomposed along the standard basis, is negative. This results in an orthant of the domain, for which the identity map is applied, whilst the remaining space is scaled by $\alpha$. The relative volume of the identity-orthant follows $2^{-n}$ for $n$-neurons in a layer, whilst the scaled region grows as $1-2^{-n}$. This is a consequence of Leaky-ReLU’s $S_{n}$ form, whilst not exhibiting a sign-flip symmetry consistent with $B_{n}$. Similar argumentation applies to any $S_{n}$-only function, though such equal and bounded orthants may be generalised. For the Leaky-ReLU case, with large widths, almost the entire domain is scaled, resulting in an exponentially vanishing volume fraction of the domain where the identity map acts. This differs from the univariate perspective, where the two domains occupy half the space.
Hence, in a multivariate view, an untrained network may be expected to have an exponentially vanishing number of representations which are not scaled. This may mean that the activation function approaches nearly linear behaviour, all scaled by $\alpha$. This may be undesirable, as representations crossing from one region to another may be beneficial for the non-linearities to affect the computation. Consequently, training may diagonalise activations to cross this boundary.
This introduces an explainable algebraic mode through which symmetries can affect representations, since the orthant-wise behaviour is characteristic of permutation symmetries. However, crucially, this is only applicable to Leaky-ReLU and fails to provide a mode through which one can explain the structure in Tanh-like networks, though it may be generalised. Hence, the overall maximal symmetry remains the summary and most effective primary causal mechanism, which may act through various modes dependent upon the function [^5] An attempt was made to produce a Leaky-ReLU-like function which partitions the space into orthants with alternating scaling and identity maps. This was termed a global-parity Leaky-ReLU and obeyed a permutation with even sign flips (denoted $D_{n}$, not to be confused with dihedral groups). It was constructed by decomposing the vector along the standard basis, then taking the sign of the product of (the sign of) components to determine either a scaling or identity map. This equal partitioning was hoped to be beneficial; however, this function failed to train in all cases and produced no influence on the representations. Perhaps future work could improve this approach by creating a working activation function with such symmetry. One could then explore its effect on representations, and explore the resultant taxonomic chain: $S_{n}\subset D_{n}\subset B_{n}\subset\mathrm{O}\left(n\right)$..
@@
Figure 21: Shows the same networks and in the same order as Fig. 19, but results arise from the $1^{\text{st}}$ hidden layer. Again, as is typical for isotropy, the bottom five rows demonstrate smooth, continuous-like representations that fill the space. In comparison, anisotropic examples in the top five rows produce narrow, ray-like distributions. Notably, these are more axis-aligned than Fig. 21, but continue to frequently display discrete Superposition-like clusters, particularly in anisotropy’s row 5.
@@ B.5 MNIST Results
Figure 22: Displays the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with $3$-hidden-layer blocks consisting of $18$ neurons per block. This architecture is comparable to those in App. D in every way except for the dimensions of CIFAR, which is exchanged for MNIST shape on input-output layers. Overall, one can observe that isotropic networks in the bottom five rows continue to produce smooth distributions that fill the representation space. In comparison, anisotropic networks on the top five rows tend to have many representations centred around the origin, and some produce axis-aligned features.
Figure 23: Displays the exact same plot as Fig. 22, with the colour map levels adjusted to make the faint representations more observable. In this plot, the axis-aligned representations of anisotropic networks are more observable. This demonstrates how anisotropic tanh produces task-agnostic structure in MNIST reconstruction autoencoders as well.
Several of the $0^{\text{th}}$ layer plots show a closer grouping of representations about the origin. This is believed to be due to many of the MNIST examples’ components reaching the extreme values for the space, 00 and $1$, with fewer intermediate values. Often, this occurs repeatedly in specific patches due to the overlapping morphologies of MNIST numbers, which may produce clustered representations inherent to the dataset. This makes the MNIST dataset more anisotropic, featuring already dense, discrete clusters about the corners of an embedded hyper-cube — significantly different from the more intermediate RGB values arising from CIFAR. Slight band-groupings of representations are also sometimes observable, likely due to the same representation being projected in multiple planes, as well as the precision of the dataset, which produces slight clusterings. However, MNIST’s greater inherent anisotropy appears not to significantly alter the results from CIFAR, likely due to the layer bottleneck embedding these higher-dimensional representations more evenly in the latent layer during initialisation.
@@
Figure 24: Displays the PPP method applied to the $1^{\text{st}}$ latent layer of an autoencoder with $3$-hidden-layer blocks consisting of $18$ neurons per block. Row-wise, these are the exact same networks as displayed in Fig. 22 and 23, but for the next latent layer. One can observe an axis-aligned structure in the top five rows, corresponding to anisotropic standard-tanh, and this is consistent with all previous experiments. In contrast, a smoother, approximately continuous distribution is observed for the bottom five rows, which are isotropic, and is consistent with this trend.
@@ Appendix C Analysis of Performance and Pathological Indications
This section displays the reconstruction errors for several of the networks tested. Normalised and Isotropic networks tend to outperform all other networks, though several exceptions do occur. The results should be taken as preliminary, since there were fewer CIFAR reconstructions computed using unnormalised datasets, due to limited compute resources and performance not being the primary study of this paper — particularly since Isotropic-Tanh is considered a placeholder activation function which is likely suboptimal and only made to appear superficially similar to standard-Tanh [2]. Nevertheless, they are indicative of a general trend where isotropic networks tend to outperform anisotropic networks on this specific task.
@@
Each plot shown represents a specific architecture, with isotropic and anisotropic activations plotted in different colours. Sometimes with their normalised and unnormalised counterparts overlayed where relevant. There are five samples per variant (each plotted faintly), along with the mean and standard deviation. These are all for the unseen testing-set reconstruction error against the number of epochs trained.
@@
Figure 26: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of $18$ neurons, and their depth increases to the right. In the leftmost plot, unnormalised isotropy has the lowest error by a very slight margin. The central two plots show that normalised isotropic networks have the lowest error by a large margin. The rightmost plot again shows unnormalised isotropy with the lowest error and a larger margin. Overall, a trend is not apparent, but isotropy tends to outperform by a greater margin.
Notably, in every example, the plots for unnormalised anisotropic networks featured significantly greater noise in reconstruction error on the testing set than the unnormalised isotropic networks. This may be indicative of smoother training dynamics, indicated for isotropic functions [2]. However, this observation is insufficiently rigorous to draw a conclusive statement regarding smoother optimisation dynamics in isotropic networks.
The tendency for unnormalised anisotropic networks to outperform normalised anisotropic networks may be due to the unnormalised dataset displaying a greater degree of initial anisotropy. This is because the unnormalised samples are bound between $\left[0,1\right]^{32\times 32\times 3}$ and rather densely cover this space. Hence, the unnormalised dataset is strongly anisotropic, only populating these directions from the origin. This may be better leveraged by the anisotropic primitives, and hence demonstrate improved reconstruction error. This is in contrast to the normalised dataset which is more uniformly distributed about the origin, as shown in App. A.2. This may also explain why isotropic networks perform better on these normalised cases. However, this does not appear to explain the later Leaky-ReLU results of Fig. 30, where anisotropic Leaky-ReLU performed more poorly on the unnormalised case compared to the normalised case — though this may be because the normalised inputs resulted in initialised activation which better straddle the non-linear primitive’s piecewise boundary.
An instance of $32$ neurons was also computed, and this example did show unnormalised-anisotropic functions performing better as shown in Fig. 27. However, these were not the best-performing models overall.
Figure 27: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction error on test-set CIFAR against the number of training epochs. The networks have a width of $32$ neurons, and $2$ hidden layer blocks. In this counterexample, the unnormalised-anisotropic network had the lowest reconstruction error.
@@
Figure 28: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) reconstruction errors for test-set MNIST against the number of training epochs, network width increases rightward over columns, and depth increases downward over rows (flipped ordering from Fig. 25). It can be observed that reconstruction error is lower in wider and less-deep networks. Isotropic networks outperform anisotropic networks in every case, except for the $3$-hidden-layer block and $32$ neurons, where anisotropic networks exhibit a slight marginal success. The best performing network is the isotropic network with 00 hidden layer blocks and $32$ neurons — this network is also thought to have the greatest representation capacity.
@@
Figure 29: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction errors on test-set MNIST against the number of training epochs. The networks have a width of $18$ neurons, and a depth of 00 hidden layer blocks. It can be observed that all networks tend to a similar reconstruction error, except for the normalised-isotropy graph, which reaches a significantly lower reconstruction error.
@@
Figure 30: These plots show the Leaky-ReLU-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of $18$ neurons, and their depth increases to the right. In all instances, normalised or unnormalised isotropic-Leaky-ReLU reaches the lowest reconstruction error by a large margin. Anisotropic networks also appear to exhibit significantly greater noise and sample-to-sample variation in test-set accuracies across the differing training epochs.
@@
These improvements may not extend to larger models such as transformers. For example, the structure of these models may also be partially a product of selection over anisotropic primitives, and may constitute a circular reinforcement of such forms [2]. Therefore, these improvements for isotropic primitives may not be applicable. A careful approach, and perhaps principled redesign/rediscovery, of larger models may be necessary for determining isotropy improvements. Nevertheless, these preliminary results offer some support for their adoption in autoencoding reconstruction tasks, even for these illustrative placeholder activation functions, which were not intended for actual usage [2].

### v2

@@ B.3 Dependence on Latent Layer of Study
Figure 14: This network shows the PPP method applied over networks of 333 hidden layer block, 181818 neurons and using input-output normalisation as described in Apps. A and D. PPP was applied to the $0^{\text{th}}$ hidden layer. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. One can observe that representations are absent in anisotropic rows 1, 2, and 4, likely a result of anti-alignment, whereas some axis-aligned structure is present in rows 3 and 5. Isotropic rows demonstrate larger, more continuous representational structures, but also demonstrate clustering such that they are not distributed smoothly over the representation space.
@@
Overall, in many networks tested, anisotropic structure appeared progressively clearer and denser in later latent layers, whilst isotropic networks tended to produce progressively more continuous distributions in later latent layers. This seemed to be a trend across many of the networks studied, but sporadic network-to-network differences also occurred. This may be due to the increased number of an/isotropies in the forward pass or indirectly through the backwards pass, yet further study may elucidate the exact reason why progressively clearer structure may emerge in later layers.
@@ B.4 Results for Leaky-ReLU
In this section, the standard Leaky-ReLU, shown in Eqn. 6, is compared with its isotropic counterpart, shown in Eqn. 7 Bird [2025a].
@@
The permutation symmetry of standard Leaky-ReLU subtly differs from that of standard tanh. Standard tanh is equivariant under both permutation and component sign-flips, termed a hyperoctahedral symmetry ($B_{n}$), whereas standard Leaky-ReLU is only equivariant under permutation ($S_{n}$). Differences in quantised structure may be indicative of such differences between symmetry groups obeyed by the functional forms.
@@
Figure 18: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 333 hidden layer blocks and width of 181818 neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. All five of the top five rows demonstrate clear anti-aligned quantisation in representations, predominantly in a single anti-aligned axis. These emerge early in training and become progressively more (approximately) discrete. This contrasts with the isotropic Leaky-ReLU examples, which approach a more continuous, evenly distributed embedding distribution, consistent with previous observations. Additionally, at relatively smaller magnitudes, the Leaky-ReLU results consistently show some deviation from the discrete-like rays; this remains unexplained.
Figure 19: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 222 hidden layer blocks and width of 181818 neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. These results are highly consistent with Fig. 18, yet the 3rd row also shows the addition of axis-aligned structure.
@@
In particular, the $0^{\text{th}}$ layer standard Leaky-ReLU results always produced two narrow beams arranged antipodally, along a single diagonal. This anti-aligned structure occurs in most cases, and the standard Leaky-ReLU only rarely displays an axis-aligned structure. This single-diagonal may be indicative of the $S_{n}$ group, which does not include sign flips, unlike tanh’s $B_{n}$ group. Therefore, only a single diagonal may appear populated, as opposed to other diagonal directions, which could result from the various mirrors applicable in the $B_{n}$ group. This nuanced difference between demonstrations appears to closely follow from the differences between symmetry groups. These differing representations can be considered to support the influence of different symmetry groups on representations, with nuances following in a taxonomic fashion from the group structures. Particularly, the orthants containing ±$\pm\vec{1}$ are the only non-degenerate orthants produced by the $S_{n}$ symmetry, and the representations appear to be aligning with these, suggesting the remaining orthants appear not to produce computationally useful maps.
@@
Particularly, Leaky-ReLU effectively uses the condition that the activation is scaled by $\alpha$ if any component of the vector, decomposed along the standard basis, is negative. This results in an orthant of the domain, for which the identity map is applied, whilst the remaining space is scaled by $\alpha$ along various combinations of components. The relative volume of the identity-orthant follows $2^{-n}$ for $n$-neurons in a layer, whilst the various scaled regions grow as $1-2^{-n}$. This is a consequence of Leaky-ReLU’s $S_{n}$ form, whilst not exhibiting a sign-flip symmetry consistent with $B_{n}$. Similar argumentation applies to any $S_{n}$-only function. For the Leaky-ReLU case, with large widths, almost the entire domain is scaled, resulting in an exponentially vanishing volume fraction of the domain where the identity map acts. This differs from the univariate perspective, where the two domains occupy half the space.
Hence, in a multivariate view, an untrained network may be expected to have an exponentially vanishing number of representations which are not scaled. This may mean that the activation function approaches a state where every representation has been scaled in some combination of components by $\alpha$. This appears to be undesirable to the network; consequently, training diagonalises it over the only orthant containing an identity map, such that activations cross this boundary and get scaled significantly differently.
This introduces an explainable algebraic mode through which symmetries can affect representations, since the orthant-wise behaviour is characteristic of permutation symmetries. However, crucially, this is a specific form of orthant effect that is only applicable to Leaky-ReLU and fails to provide a mode through which one can explain the structure in Tanh-like networks, though it may be generalised. The generalised orthant explanation continues to hold. Hence, the overall maximal symmetry remains the summary and most effective primary causal mechanism, which may act through various modes dependent upon the function [^8] An attempt was made to produce a Leaky-ReLU-like function which partitions the space into orthants with alternating scaling and identity maps. This was termed a global-parity Leaky-ReLU and obeyed a permutation with even sign flips (denoted $D_{n}$, not to be confused with dihedral groups). It was constructed by decomposing the vector along the standard basis, then taking the sign of the product of (the sign of) components to determine either a scaling or identity map. This equal partitioning was hoped to be beneficial; however, this function failed to train in all cases and produced no influence on the representations. Perhaps future work could improve this approach by creating a working activation function with such symmetry. One could then explore its effect on representations, and explore the resultant taxonomic chain: $S_{n}\subset D_{n}\subset B_{n}\subset\mathrm{O}\left(n\right)$..
@@
Figure 21: Shows the same networks and in the same order as Fig. 19, but results arise from the $1^{\text{st}}$ hidden layer. Again, as is typical for isotropy, the bottom five rows demonstrate smooth, continuous-like representations that fill the space. In comparison, anisotropic examples in the top five rows produce narrow, ray-like distributions. Notably, these are more axis-aligned than Fig. 21, but continue to frequently display discrete Superposition-like clusters, particularly in anisotropy’s row 5. Unlike the prior plots, the anisotropic networks all align upwards and towards the left, potentially indicating a differing desirable orthant mapping in this case.
@@ B.5 MNIST Results
Figure 22: Displays the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 333-hidden-layer blocks consisting of 181818 neurons per block. This architecture is comparable to those in App. D in every way except for the dimensions of CIFAR, which is exchanged for MNIST shape on input-output layers. Overall, one can observe that isotropic networks in the bottom five rows continue to produce smooth distributions that fill the representation space. In comparison, anisotropic networks on the top five rows tend to have many representations centred around the origin, and some produce axis-aligned features.
Figure 23: Displays the exact same plot as Fig. 22, with the colour map levels adjusted in post-processing to make the faint representations more observable. In this plot, the axis-aligned representations of anisotropic networks are more observable. This demonstrates how anisotropic tanh produces task-agnostic structure in MNIST reconstruction autoencoders as well.
Several of the $0^{\text{th}}$ layer plots show a closer grouping of representations about the origin. This is believed to be due to many of the MNIST examples’ components reaching the extreme values for the space, 0 and 111, with fewer intermediate values. Often, this occurs repeatedly in specific patches due to the overlapping morphologies of MNIST numbers, which may produce clustered representations inherent to the dataset. This makes the MNIST dataset more anisotropic, featuring already dense, discrete clusters about the corners of an embedded hyper-cube — significantly different from the more intermediate RGB values arising from CIFAR. Slight band-groupings of representations are also sometimes observable, likely due to the same representation being projected in multiple planes, as well as the precision of the dataset, which produces slight clusterings. However, MNIST’s greater inherent anisotropy appears not to significantly alter the results from CIFAR, likely due to the layer bottleneck embedding these higher-dimensional representations more evenly in the latent layer during initialisation.
@@
Figure 24: Displays the PPP method applied to the $1^{\text{st}}$ latent layer of an autoencoder with 333-hidden-layer blocks consisting of 181818 neurons per block. Row-wise, these are the exact same networks as displayed in Fig. 22 and 23, but for the next latent layer. One can observe an axis-aligned structure in the top five rows, corresponding to anisotropic standard-tanh, and this is consistent with all previous experiments. In contrast, a smoother, approximately continuous distribution is observed for the bottom five rows, which are isotropic, and is consistent with this trend.
@@ Appendix C Analysis of Performance and Pathological Indications
This section displays the reconstruction errors for several of the networks tested. Normalised and Isotropic networks tend to outperform all other networks, though several exceptions do occur. The results should be taken as preliminary, since there were fewer CIFAR reconstructions computed using unnormalised datasets, due to limited compute resources and performance not being the primary study of this paper — particularly since Isotropic-Tanh is considered a placeholder activation function which is likely suboptimal and only made to appear superficially similar to standard-Tanh Bird [2025a]. Nevertheless, they are indicative of a general trend where isotropic networks tend to outperform anisotropic networks on this specific task.
@@
Each plot shown represents a specific architecture, with isotropic and anisotropic activations plotted in different colours. Sometimes with their normalised and unnormalised counterparts overlaid where relevant. There are five samples per variant (each plotted faintly), along with the mean and standard deviation. These are all for the unseen testing-set reconstruction error against the number of epochs trained.
@@
Figure 26: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of 181818 neurons, and their depth increases to the right. In the leftmost plot, unnormalised isotropy has the lowest error by a very slight margin. The central two plots show that normalised isotropic networks have the lowest error by a large margin. The rightmost plot again shows unnormalised isotropy with the lowest error and a larger margin. Overall, a trend is not apparent, but isotropy tends to outperform by a greater margin.
Notably, in every example, the plots for unnormalised anisotropic networks featured significantly greater noise in reconstruction error on the testing set than the unnormalised isotropic networks. This may be indicative of smoother training dynamics, indicated for isotropic functions Bird [2025a]. However, this observation is insufficiently rigorous to draw a conclusive statement regarding smoother optimisation dynamics in isotropic networks.
The tendency for unnormalised anisotropic networks to outperform normalised anisotropic networks may be due to the unnormalised dataset displaying a greater degree of initial anisotropy. This is because the unnormalised samples are bound between $\left[0,1\right]^{32\times 32\times 3}$ and rather densely cover this space. Hence, the unnormalised dataset is strongly anisotropic, only populating these directions from the origin. This may be better leveraged by the anisotropic primitives, and hence demonstrate improved reconstruction error. This is in contrast to the normalised dataset, which is more uniformly distributed about the origin, as shown in App. A.2. This may also explain why isotropic networks perform better on these normalised cases. However, this does not appear to explain the later Leaky-ReLU results of Fig. 30, where anisotropic Leaky-ReLU performed more poorly on the unnormalised case compared to the normalised case — though this may be because the normalised inputs resulted in initialised activation which better straddle the non-linear primitive’s piecewise boundary.
An instance of 323232 neurons was also computed, and this example did show unnormalised-anisotropic functions performing better as shown in Fig. 27. However, these were not the best-performing models overall.
Figure 27: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction error on test-set CIFAR against the number of training epochs. The networks have a width of 323232 neurons, and 222 hidden layer blocks. In this counterexample, the unnormalised-anisotropic network had the lowest reconstruction error.
@@
Figure 28: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) reconstruction errors for test-set MNIST against the number of training epochs, network width increases rightward over columns, and depth increases downward over rows (flipped ordering from Fig. 25). It can be observed that reconstruction error is lower in wider and less-deep networks. Isotropic networks outperform anisotropic networks in every case, except for the 333-hidden-layer block and 323232 neurons, where anisotropic networks exhibit a slight marginal success. The best performing network is the isotropic network with 0 hidden layer blocks and 323232 neurons — this network is also thought to have the greatest representation capacity.
@@
Figure 29: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction errors on test-set MNIST against the number of training epochs. The networks have a width of 181818 neurons, and a depth of 0 hidden layer blocks. It can be observed that all networks tend to a similar reconstruction error, except for the normalised-isotropy graph, which reaches a significantly lower reconstruction error.
@@
Figure 30: These plots show the Leaky-ReLU-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of 181818 neurons, and their depth increases to the right. In all instances, normalised or unnormalised isotropic-Leaky-ReLU reaches the lowest reconstruction error by a large margin. Anisotropic networks also appear to exhibit significantly greater noise and sample-to-sample variation in test-set accuracies across the differing training epochs.
@@
These improvements may not extend to larger models such as transformers. For example, the structure of these models may also be partially a product of selection over anisotropic primitives, and may constitute a circular reinforcement of such forms Bird [2025a]. Therefore, these improvements for isotropic primitives may not be applicable. A careful approach, and perhaps symmetry-principled redesign/rediscovery, of larger models may be necessary for determining isotropy improvements. Nevertheless, these preliminary results offer some support for their adoption in autoencoding reconstruction tasks, even for these illustrative placeholder activation functions, which were not intended for actual usage Bird [2025a].